OpenStudy (anonymous):

find the second derivative of 2x^2-y^2=1 using implicit differentiation

OpenStudy (anonymous):

start with \[4x-2yy'=0\] then solve for \(y'\)

OpenStudy (anonymous):

Okay that gave me 2x/y and that's where I get lost

OpenStudy (anonymous):

looks good to me

OpenStudy (anonymous):

unless you solve for \(y\) explicitly in terms of \(x\) your answer will have both an \(x\) and a \(y\) in it

OpenStudy (anonymous):

I understand that I 'm trying to find teh second derivative.

OpenStudy (anonymous):

oooh second derivative!! my fault, sorry now you have to take the derivative of \(\frac{2x}{y}\) using the quotient rule

OpenStudy (anonymous):

\[\frac{2y-2xy'}{y^2}\] is the first step via the quotient rule

OpenStudy (anonymous):

I got that, I just don't know where to go from there!

OpenStudy (anonymous):

second step is to replace \(y'\) by \(\frac{2x}{y}\) and then some algebra

OpenStudy (anonymous):

\[\frac{2y-2xy'}{y^2}\] \[\frac{2y-2x\frac{2x}{y}}{y^2}\]

OpenStudy (anonymous):

so, 2y-(2x^2/y)/y^2

OpenStudy (anonymous):

multiply top and bottom by \(y\) to clear the fraction, get \[\frac{2y^2-4x^2}{y^3}\]

OpenStudy (anonymous):

OH!!

OpenStudy (anonymous):

yeah you are right, need to clean it up a little that's all

OpenStudy (anonymous):

we can clean it up even more

OpenStudy (anonymous):

\[2y^2-4x^2=-2(2x^2-y^2)=-2\times 1\] pretty tricky huh?

OpenStudy (anonymous):

Okay, thanks for the quick response and help!! My answer key doesn't want the simplified version, but neat trick!

OpenStudy (anonymous):

yw